OpenStudy (anonymous):

I feel like I'm missing something here. Someone point me in the right direction? Is there a specific formula to create a system of equations? Create a system of equations that includes one linear equation and one quadratic equation. Part 1. Show all work to solving your system of equations algebraically. Part 2. Graph your system of equations and show the solution graphically to verify your solution.

OpenStudy (loser66):

 how about this?

OpenStudy (loser66):

you can arbitrary write down the equation of the line, let make it as simple as possible, y =x and the equation of the parabola y = x^2 then, your system is \[y =x^2\\y=x\] and the question is "Solve"

OpenStudy (loser66):

and how to solve is your duty. It is not hard, right?